# Reality conditions for the KdV equation and quasi-periodic solutions in finite phase spaces

### v2

@@ 4.1. Spectral curve
(26) $-y^{2}+H(z)=0.$
Indeed, the spectral curve of each hamiltonian system in the hierarchy is defined by the characteristic polynomial of $\Psi$, namely $\det\big{(}\Psi(z)-y\big{)}=0$. Recall, that $H$ is the polynomial (15) of degree $12N+$ 1. All coefficients are integrals of motion: $h_{0}$, …, $h_{N-1}$ serve as hamiltonians, and $h_{N}$, …, $h_{2N}$ introduce constraints (17), which fix an orbit, and $h_{2N+1}=\mathrm{b}^{2}$.
@@ 4.2. Canonical coordinates
(34) $\textsf{B}_{\text{h}}=\begin{pmatrix}0&0&\beta_{1}&\beta_{3}&\dots&\beta_{2N-3}&\beta_{2N-1}\\ 0&0&0&\beta_{1}&\beta_{3}&\dots&\beta_{2N-3}\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ 0&0&0&0&\dots&\beta_{1}&\beta_{3}\\ 0&0&0&0&\dots&0&\beta_{1}\end{pmatrix},$
(45) $\textsf{A}_{\text{h}}=\begin{pmatrix}\sum_{n=1}^{N}\alpha_{2(N-n)}\alpha_{2(n-1)}\\ \sum_{n=2}^{N}\alpha_{2(N-n)}\alpha_{2(n-2)}\\ \vdots\\ 2\alpha_{0}\alpha_{2}\\ \alpha_{0}^{2}\end{pmatrix},\qquad\bm{h}=\begin{pmatrix}h_{N-1}\\ h_{N-2}\\ \vdots\\ h_{1}\\ h_{0}\end{pmatrix}.$
@@
On the other hand, hamiltonians $h_{0}$, …, $h_{N-1}$ can be found from the equation of the spectral curve, taken at $N$ points which form a non-special [^4] A positive divisor of degree $g$ on a hyperelliptic curve of genus $g$ contains no pair of points in involution. divisor. Namely, with $i=1$, …, $N$
$-y_{i}^{2}+\mathrm{b}^{2}z_{i}^{2N+1}+c_{2N}z_{i}^{2N}+\cdots+c_{N}z_{i}^{N}+h_{N-1}z_{i}^{N-1}+\dots+h_{1}z_{i}+h_{0}=0,$
@@
$-\bm{y}+Z_{\text{c}}\bm{c}+Z_{\text{h}}\bm{h}=0,$
@@
$Z_{\text{c}}=\begin{pmatrix}z_{1}^{2N+1}&z_{1}^{2N}&\dots&z_{1}^{N}\\ z_{2}^{2N+1}&z_{2}^{2N}&\dots&z_{2}^{N}\\ \vdots&\vdots&\ddots&\vdots\\ z_{N}^{2N+1}&z_{N}^{2N}&\dots&z_{N}^{N}\end{pmatrix},\quad Z_{\text{h}}=\begin{pmatrix}z_{1}^{N-1}&\dots&z_{1}&1\\ z_{2}^{N-1}&\dots&z_{2}&1\\ \vdots&\ddots&\vdots&\vdots\\ z_{N}^{N-1}&\dots&z_{N}&1\end{pmatrix},\quad\bm{y}=\begin{pmatrix}y_{1}^{2}\\ y_{2}^{2}\\ \vdots\\ y_{N}^{2}\end{pmatrix}.$
The matrix $Z_{\text{h}}$ is square and invertible. Thus,
(47) $\bm{h}=Z_{\text{h}}^{-1}\big{(}\bm{y}-Z_{\text{c}}\bm{c}\big{)}.$
@@
$\textsf{B}_{\text{h}}\textsf{B}_{\text{c}}^{-1}(\bm{c}-\textsf{A}_{\text{c}})+\textsf{A}_{\text{h}}=Z_{\text{h}}^{-1}\big{(}\bm{y}-Z_{\text{c}}\bm{c}\big{)}.$
@@
(48a) $\textsf{B}_{\text{h}}\textsf{B}_{\text{c}}^{-1}=-Z_{\text{h}}^{-1}Z_{\text{c}},$
(48b) $-\textsf{B}_{\text{h}}\textsf{B}_{\text{c}}^{-1}\textsf{A}_{\text{c}}+\textsf{A}_{\text{h}}=Z_{\text{h}}^{-1}\bm{y}.$
@@
$Z_{\text{h}}\textsf{B}_{\text{h}}+Z_{\text{c}}\textsf{B}_{\text{c}}=0,$
which is equivalent to $\beta(z_{i})=0$, since $Z_{\text{h}}$ is the Vandermonde matrix. Then from (48b) we obtain
$Z_{\text{c}}\textsf{A}_{\text{c}}+Z_{\text{h}}\textsf{A}_{\text{h}}=\bm{y},$
which is equivalent to $y_{i}^{2}-\alpha(z_{i})^{2}=0$. Thus, points $(z_{i},y_{i})$ are defined by
$\beta(z_{i})=0,\qquad y_{i}^{2}-\alpha(z_{i})^{2}=0,\quad i=1,\dots,N.$
@@
Suppose an orbit $\mathcal{O}\subset\mathcal{M}_{N}^{\circ}$ has the coordinates $(\beta_{2m+1},\alpha_{2m})$, $m=0$, …, $N-1$, as above. Then the new coordinates $(z_{i},y_{i})$, $i=1$, …, $N$, defined by the formulas
(49) $\beta(z_{i})=0,\qquad y_{i}=\epsilon\alpha(z_{i}),\quad i=1,\dots,N,$
@@
  1. (1)
    a pair $(z_{i},y_{i})$ is a root of the characteristic polynomial (26).
  2. (2)
    a pair $(z_{i},y_{i})$ is canonically conjugate with respect to the Lie-Poisson bracket (13):
    (50) $\{z_{i},z_{j}\}=0,\qquad\{z_{i},y_{j}\}=-\epsilon\,\delta_{i,j},\qquad\{y_{i},y_{j}\}=0.$
  3. (3)
@@
$\{z_{i},y_{j}\}=\sum_{n+m=0}^{N-1}\bigg{(}\frac{\partial z_{i}}{\beta_{2m+1}}\frac{\partial y_{j}}{\alpha_{2n}}-\frac{\partial z_{i}}{\alpha_{2n}}\frac{\partial y_{j}}{\beta_{2m+1}}\bigg{)}\{\beta_{2m+1},\alpha_{2n}\}\\ =\frac{-\epsilon}{\beta^{\prime}(z_{i})}\sum_{m+n=0}^{N-1}z_{i}^{m}z_{j}^{n}\beta_{2(m+n)+3}=\frac{-\epsilon}{\beta^{\prime}(z_{i})}\frac{\beta(z_{i})-\beta(z_{j})}{z_{i}-z_{j}},$
@@
$\frac{\partial z_{i}}{\beta_{2m+1}}=-\frac{z_{i}^{m}}{\beta^{\prime}(z_{i})},\qquad\frac{\partial z_{i}}{\alpha_{2n}}=0,\qquad\frac{\partial y_{i}}{\alpha_{2n}}=\epsilon z_{i}^{n},\qquad\frac{\partial y_{i}}{\beta_{2m+1}}=-\epsilon z_{i}^{m}\frac{\alpha^{\prime}(z_{i})}{\beta^{\prime}(z_{i})}.$
As $i\neq j$, it is evident that $\{z_{i},y_{j}\}=0$, due to $\beta(z_{i})=\beta(z_{j})=0$. As $i=j$, we get
$\{z_{i},y_{i}\}=\lim_{z_{j}\to z_{i}}\frac{-\epsilon}{\beta^{\prime}(z_{i})}\frac{\beta(z_{i})-\beta(z_{j})}{z_{i}-z_{j}}=-\epsilon.$
@@
$\{y_{i},y_{j}\}=\sum_{n+m=0}^{N-1}\bigg{(}\frac{\partial y_{i}}{\beta_{2m+1}}\frac{\partial y_{j}}{\alpha_{2n}}-\frac{\partial y_{i}}{\alpha_{2n}}\frac{\partial y_{j}}{\beta_{2m+1}}\bigg{)}\{\beta_{2m+1},\alpha_{2n}\}\\ =-\epsilon^{2}\sum_{m+n=0}^{N-1}\bigg{(}z_{i}^{m}z_{j}^{n}\frac{\alpha^{\prime}(z_{i})}{\beta^{\prime}(z_{i})}-z_{i}^{n}z_{j}^{m}\frac{\alpha^{\prime}(z_{j})}{\beta^{\prime}(z_{j})}\bigg{)}\beta_{2(m+n)+3}\\ =-\epsilon^{2}\frac{\beta(z_{i})-\beta(z_{j})}{z_{i}-z_{j}}\bigg{(}\frac{\alpha^{\prime}(z_{i})}{\beta^{\prime}(z_{i})}-\frac{\alpha^{\prime}(z_{j})}{\beta^{\prime}(z_{j})}\bigg{)}.$
Thus, $\{y_{i},y_{j}\}=0$, due to $\beta(z_{i})=\beta(z_{j})=0$.
@@ 5.1. Uniformization of the spectral curve
After separation of variables, we came to the Jacobi inversion problem for a non-special divisor of $N$ points $\{(z_{i},y_{i})\}_{i=1}^{N}$ on a hyperelliptic curve of genus $N$
(52) $f(z,y)\equiv-y^{2}+\mathrm{b}^{2}z^{2N+1}+c_{2N}z^{2N}+\cdots+c_{N}z^{N}\\ +h_{N-1}z^{N-1}+\dots+h_{1}z+h_{0}=0.$
@@
(53a) $\mathrm{d}u_{2n-1}=\frac{\mathrm{b}z^{g-n}\mathrm{d}z}{\partial_{y}f(z,y)},\quad n=1,\dots,g,$
(53b) $\mathrm{d}r_{2n-1}=\frac{\mathrm{d}z}{\mathrm{b}\partial_{y}f(z,y)}\sum_{j=1}^{2n-1}(2n-j)h_{2g+2-j}z^{g+n-j},$
which follow from (3a), (3b), after reducing (52) to the form (2) by applying the transformation $f(z,y)\mapsto f(z,\mathrm{b}y)/\mathrm{b}^{2}$.
Solution of the the Jacobi inversion problem on a hyperelliptic curve can be found in [3, §216], see also [7, Theorem 2.2]. On the curve (52), the Abel pre-image of $\mathcal{A}(D)=u\in\operatorname{\mathfrak{J}}$, which is a non-special positive divisor $D=\sum_{k=1}^{N}(z_{k},y_{k})$, is obtained from the system
@@
(54b) $\frac{2y}{\mathrm{b}}+\sum_{k=1}^{N}z^{N-k}\wp_{1,1,2k-1}(u)=0.$
According to (49), the $N$ values $z_{i}$ are zeros of the polynomial $\beta(z)$, and the $N$ values $y_{i}$ are obtained from $y_{i}=-\alpha(z_{i})$. Thus,
@@
This solution arose in [7, Theorem 4.12]
@@
which holds for hyperelliptic Weierstrass-$\wp$-functions in any genus. The relation corresponds to a curve of the form (2). Assigning $\lambda_{2}=c_{2N}/\mathrm{b}^{2}$, $\lambda_{4}=c_{2N-1}/\mathrm{b}^{2}$, we get the relation for the spectral curve (52). Differentiation with respect to $u_{1}$ brings (57) to the KdV equaiton (22) in terms of Weierstrass-$\wp$-functions.
@@ 5.2. Equation of motion in variables of separation
$\frac{\mathrm{d}}{\mathrm{d}x}\beta(z)=2\mathrm{a}\beta(z)-2\mathrm{b}\alpha(z),$
$\frac{\mathrm{d}}{\mathrm{d}t}\beta(z)=2(\mathrm{a}z+\alpha_{2N-2})\beta(z)-2(\mathrm{b}z+\beta_{2N-1})\alpha(z),$
where all dynamic variables are functions of $x$ and $t$. Therefore, zeros of $\beta(z)$ are functions of $x$ and $t$ as well, namely $\beta(z)=\mathrm{b}\prod_{k=1}^{N}(z-z_{k}(x,t))$. Then
$\frac{\mathrm{d}}{\mathrm{d}x}\log\beta(z)=-\frac{1}{z-z_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}x}=2\mathrm{a}-2\mathrm{b}\frac{\alpha(z)}{\beta(z)},\quad k=1,\dots,N,$
$\frac{\mathrm{d}}{\mathrm{d}t}\log\beta(z)=-\frac{1}{z-z_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}t}=2(\mathrm{a}z+\alpha_{2N-2})-2(\mathrm{b}z+\beta_{2N-1})\frac{\alpha(z)}{\beta(z)}.$
@@
$\frac{\mathrm{d}z_{k}}{\mathrm{d}x}=\frac{2y_{k}}{\prod_{j\neq k}^{N}(z_{k}-z_{j})},$ $\frac{\mathrm{d}z_{k}}{\mathrm{d}t}=-\frac{2y_{k}\sum_{j\neq k}z_{j}}{\prod_{j\neq k}(z_{k}-z_{j})}.$
Now, let $D$ be a divisor of points $\{(z_{k},y_{k})\}_{k=1}^{N}$ defined by (49). Abel’s map
$u=\mathcal{A}(D)=\sum_{k=1}^{N}\int_{\infty}^{(z_{k},y_{k})}\mathrm{d}u=\sum_{k=1}^{N}\int_{\infty}^{(z_{k},y_{k})}\begin{pmatrix}1\\ z\\ \vdots\\ z^{N-1}\end{pmatrix}\frac{\mathrm{b}\mathrm{d}z}{-2y}$
depends on $x$ and $t$, since the points $(z_{k},y_{k})$ are functions of $x$ and $t$. Then
$\frac{\mathrm{d}u_{2n-1}}{\mathrm{d}x}=\sum_{k=1}^{N}\frac{\mathrm{b}z_{k}^{N-n}}{-2y_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}x}=\sum_{k=1}^{N}\frac{-\mathrm{b}z_{k}^{N-n}}{\prod_{j\neq k}^{N}(z_{k}-z_{j})}=-\mathrm{b}\delta_{n,1},$
$\frac{\mathrm{d}u_{2n-1}}{\mathrm{d}t}=\sum_{k=1}^{N}\frac{\mathrm{b}z_{k}^{N-n}}{-2y_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}t}=\sum_{k=1}^{N}\frac{\mathrm{b}z_{k}^{N-n}\sum_{j\neq k}z_{j}}{\prod_{j\neq k}(z_{k}-z_{j})}=-\mathrm{b}\delta_{n,2}.$
Thus, $u_{1}=-\mathrm{b}x+C_{1}$, $u_{3}=-\mathrm{b}t+C_{3}$, and $u_{2n-1}=C_{2n-1}$, $n=3$, …, $N$, where $C_{2n-1}$ are constant.
@@
(61) $\beta(x,t)=-\mathrm{b}\wp_{1,1}(-\mathrm{b}x+C_{1},-\mathrm{b}t+C_{3},C_{5},\dots,C_{2N-1}).$
Since $\mathrm{b}^{2}\in\operatorname{\mathbb{R}}$, we have two possibilities: (i) $\mathrm{b}$ is real, or (ii) $\mathrm{b}$ is purely imaginary. In the case (i), the first two arguments $u_{1}$, $u_{3}$ of $\wp_{1,1}$ run along lines parallel to the real axes. In the case (ii), the first two arguments $u_{1}$, $u_{3}$ of $\wp_{1,1}$ run along lines parallel to the imaginary axes. If $\beta$ describes a quasi-periodic wave, none of the mentioned lines coincides with the real or imaginary axis, due to the singularity of $\wp_{1,1}$ at $u=0$.
Next, we find such constants $C_{2n-1}$ that $\beta$ in (61) is real-valued. We call this the reality conditions.

### v4

@@ 4.1. Spectral curve
(30) $-w^{2}+H(z)=0.$
Indeed, the spectral curve of each hamiltonian system in the hierarchy is defined by the characteristic polynomial of $\Psi$, namely $\det\big{(}\Psi(z)-w\big{)}=0$. Recall, that $H$ is the polynomial (19) of degree $12N+$ 1. All coefficients are integrals of motion: $h_{0}$, …, $h_{N-1}$ serve as hamiltonians, and $h_{N}$, …, $h_{2N}$ introduce constraints (21), which fix an orbit, and $h_{2N+1}=\mathrm{b}^{2}$.
@@ 4.2. Canonical coordinates
$\textsf{B}_{\text{h}}=\begin{pmatrix}0&0&\beta_{1}&\beta_{3}&\dots&\beta_{2N-3}&\beta_{2N-1}\\ 0&0&0&\beta_{1}&\beta_{3}&\dots&\beta_{2N-3}\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ 0&0&0&0&\dots&\beta_{1}&\beta_{3}\\ 0&0&0&0&\dots&0&\beta_{1}\end{pmatrix},$
$\textsf{A}_{\text{h}}=\begin{pmatrix}\sum_{n=1}^{N}\alpha_{2(N-n)}\alpha_{2(n-1)}\\ \sum_{n=2}^{N}\alpha_{2(N-n)}\alpha_{2(n-2)}\\ \vdots\\ 2\alpha_{0}\alpha_{2}\\ \alpha_{0}^{2}\end{pmatrix},\qquad\bm{h}=\begin{pmatrix}h_{N-1}\\ h_{N-2}\\ \vdots\\ h_{1}\\ h_{0}\end{pmatrix}.$
@@
On the other hand, hamiltonians $h_{0}$, …, $h_{N-1}$ can be found from the equation of the spectral curve, taken at $N$ points which form a non-special [^4] Here a non-special divisor is supposed to be a positive divisor of degree $g$ on a hyperelliptic curve of genus $g$ contains no pair of points in involution. divisor. Namely, with $i=1$, …, $N$
$-w_{i}^{2}+\mathrm{b}^{2}z_{i}^{2N+1}+c_{2N}z_{i}^{2N}+\cdots+c_{N}z_{i}^{N}+h_{N-1}z_{i}^{N-1}+\dots+h_{1}z_{i}+h_{0}=0,$
@@
$-\bm{w}+\mathrm{Z}_{\text{c}}\bm{c}+\mathrm{Z}_{\text{h}}\bm{h}=0,$
@@
$\mathrm{Z}_{\text{c}}=\begin{pmatrix}z_{1}^{2N+1}&z_{1}^{2N}&\dots&z_{1}^{N}\\ z_{2}^{2N+1}&z_{2}^{2N}&\dots&z_{2}^{N}\\ \vdots&\vdots&\ddots&\vdots\\ z_{N}^{2N+1}&z_{N}^{2N}&\dots&z_{N}^{N}\end{pmatrix},\quad\mathrm{Z}_{\text{h}}=\begin{pmatrix}z_{1}^{N-1}&\dots&z_{1}&1\\ z_{2}^{N-1}&\dots&z_{2}&1\\ \vdots&\ddots&\vdots&\vdots\\ z_{N}^{N-1}&\dots&z_{N}&1\end{pmatrix},\quad\bm{w}=\begin{pmatrix}w_{1}^{2}\\ w_{2}^{2}\\ \vdots\\ w_{N}^{2}\end{pmatrix}.$
The matrix $\mathrm{Z}_{\text{h}}$ is square and invertible. Thus,
(34) $\bm{h}=\mathrm{Z}_{\text{h}}^{-1}\big{(}\bm{w}-\mathrm{Z}_{\text{c}}\bm{c}\big{)}.$
@@
$\textsf{B}_{\text{h}}\textsf{B}_{\text{c}}^{-1}(\bm{c}-\textsf{A}_{\text{c}})+\textsf{A}_{\text{h}}=\mathrm{Z}_{\text{h}}^{-1}\big{(}\bm{w}-\mathrm{Z}_{\text{c}}\bm{c}\big{)}.$
@@
(35a) $\textsf{B}_{\text{h}}\textsf{B}_{\text{c}}^{-1}=-\mathrm{Z}_{\text{h}}^{-1}\mathrm{Z}_{\text{c}},$
(35b) $-\textsf{B}_{\text{h}}\textsf{B}_{\text{c}}^{-1}\textsf{A}_{\text{c}}+\textsf{A}_{\text{h}}=\mathrm{Z}_{\text{h}}^{-1}\bm{w}.$
@@
$\mathrm{Z}_{\text{h}}\textsf{B}_{\text{h}}+\mathrm{Z}_{\text{c}}\textsf{B}_{\text{c}}=0,$
which is equivalent to $\beta(z_{i})=0$, since $\mathrm{Z}_{\text{h}}$ is the Vandermonde matrix. Then from (35b) we obtain
$\mathrm{Z}_{\text{c}}\textsf{A}_{\text{c}}+\mathrm{Z}_{\text{h}}\textsf{A}_{\text{h}}=\bm{w},$
which is equivalent to $w_{i}^{2}-\alpha(z_{i})^{2}=0$. Thus, points $(z_{i},w_{i})$ are defined by
$\beta(z_{i})=0,\qquad w_{i}^{2}-\alpha(z_{i})^{2}=0,\quad i=1,\dots,N.$
@@
Suppose an orbit $\mathcal{O}\subset\mathcal{M}_{N}^{\circ}$ has the coordinates $(\beta_{2m+1},\alpha_{2m})$, $m=0$, …, $N-1$, as above. Then the new coordinates $(z_{i},w_{i})$, $i=1$, …, $N$, defined by the formulas
(36) $\beta(z_{i})=0,\qquad w_{i}=\epsilon\alpha(z_{i}),\quad i=1,\dots,N,$
@@
  1. 1)
    a pair $(z_{i},w_{i})$ is a root of the characteristic polynomial (30).
  2. 2)
    a pair $(z_{i},w_{i})$ is canonically conjugate with respect to the Lie-Poisson bracket (17):
    (37) $\{z_{i},z_{j}\}=0,\qquad\{z_{i},w_{j}\}=-\epsilon\,\delta_{i,j},\qquad\{w_{i},w_{j}\}=0.$
  3. 3)
@@
$\{z_{i},w_{j}\}=\sum_{n+m=0}^{N-1}\bigg{(}\frac{\partial z_{i}}{\beta_{2m+1}}\frac{\partial w_{j}}{\alpha_{2n}}-\frac{\partial z_{i}}{\alpha_{2n}}\frac{\partial w_{j}}{\beta_{2m+1}}\bigg{)}\{\beta_{2m+1},\alpha_{2n}\}\\ =\frac{-\epsilon}{\beta^{\prime}(z_{i})}\sum_{m+n=0}^{N-1}z_{i}^{m}z_{j}^{n}\beta_{2(m+n)+3}=\frac{-\epsilon}{\beta^{\prime}(z_{i})}\frac{\beta(z_{i})-\beta(z_{j})}{z_{i}-z_{j}},$
@@
$\frac{\partial z_{i}}{\beta_{2m+1}}=-\frac{z_{i}^{m}}{\beta^{\prime}(z_{i})},\qquad\frac{\partial z_{i}}{\alpha_{2n}}=0,\qquad\frac{\partial w_{i}}{\alpha_{2n}}=\epsilon z_{i}^{n},\qquad\frac{\partial w_{i}}{\beta_{2m+1}}=-\epsilon z_{i}^{m}\frac{\alpha^{\prime}(z_{i})}{\beta^{\prime}(z_{i})}.$
As $i\neq j$, it is evident that $\{z_{i},w_{j}\}=0$, due to $\beta(z_{i})=\beta(z_{j})=0$. As $i=j$, we get
$\{z_{i},w_{i}\}=\lim_{z_{j}\to z_{i}}\frac{-\epsilon}{\beta^{\prime}(z_{i})}\frac{\beta(z_{i})-\beta(z_{j})}{z_{i}-z_{j}}=-\epsilon.$
@@
$\{w_{i},w_{j}\}=\sum_{n+m=0}^{N-1}\bigg{(}\frac{\partial w_{i}}{\beta_{2m+1}}\frac{\partial w_{j}}{\alpha_{2n}}-\frac{\partial w_{i}}{\alpha_{2n}}\frac{\partial w_{j}}{\beta_{2m+1}}\bigg{)}\{\beta_{2m+1},\alpha_{2n}\}\\ =-\epsilon^{2}\sum_{m+n=0}^{N-1}\bigg{(}z_{i}^{m}z_{j}^{n}\frac{\alpha^{\prime}(z_{i})}{\beta^{\prime}(z_{i})}-z_{i}^{n}z_{j}^{m}\frac{\alpha^{\prime}(z_{j})}{\beta^{\prime}(z_{j})}\bigg{)}\beta_{2(m+n)+3}\\ =-\epsilon^{2}\frac{\beta(z_{i})-\beta(z_{j})}{z_{i}-z_{j}}\bigg{(}\frac{\alpha^{\prime}(z_{i})}{\beta^{\prime}(z_{i})}-\frac{\alpha^{\prime}(z_{j})}{\beta^{\prime}(z_{j})}\bigg{)}.$
Thus, $\{w_{i},w_{j}\}=0$, due to $\beta(z_{i})=\beta(z_{j})=0$.
@@ 5.1. Uniformization of the spectral curve
After separation of variables, we came to the Jacobi inversion problem for a non-special divisor of $N$ points $\{(z_{k},w_{k})\}_{k=1}^{N}$ on a hyperelliptic curve of genus $N$
(39) $0=F(z,w)\equiv-w^{2}+\mathrm{b}^{2}z^{2N+1}+c_{2N}z^{2N}+\cdots+c_{N}z^{N}\\ +h_{N-1}z^{N-1}+\dots+h_{1}z+h_{0}.$
@@
(40a) $\mathrm{d}u_{2n-1}=\frac{\mathrm{b}z^{N-n}\mathrm{d}z}{\partial_{w}F(z,w)},\quad n=1,\dots,g,$
(40b) $\mathrm{d}r_{2n-1}=\frac{\mathrm{d}z}{\mathrm{b}\partial_{w}F(z,w)}\sum_{j=1}^{2n-1}(2n-j)h_{2N+2-j}z^{N+n-j},$
which follow from (4a), (4b), after reducing (39) to the form (3) by applying the transformation: $z\mapsto x$, $w\mapsto\mathrm{b}y$, $F(z,w)\mapsto f(x,y)=F(x,\mathrm{b}y)/\mathrm{b}^{2}$.
On the curve (39), a solution of the Jacobi inversion problem, that is the Abel pre-image of $\mathcal{A}(D)=u\in\operatorname{\mathfrak{J}}$, with a non-special positive divisor $D=\sum_{k=1}^{N}(z_{k},w_{k})$, is given by the system
@@
(41b) $\frac{2w}{\mathrm{b}}+\sum_{k=1}^{N}z^{N-k}\wp_{1,1,2k-1}(u)=0.$
According to (36), the $N$ values $z_{i}$ are zeros of the polynomial $\beta(z)$, and the $N$ values $w_{i}$ satisfy $w_{i}=-\alpha(z_{i})$. Thus,
@@
This solution arose in [9, Theorem 4.12].
@@
which holds for hyperelliptic Weierstrass-$\wp$-functions in any genus. The relation corresponds to a curve of the form (3). Assigning $\lambda_{2}=c_{2N}/\mathrm{b}^{2}$, $\lambda_{4}=c_{2N-1}/\mathrm{b}^{2}$, we get the relation for the spectral curve (39). Differentiation with respect to $u_{1}$ transforms (44) into the KdV equaiton (26) in terms of Weierstrass-$\wp$-functions.
@@ 5.2. Equation of motion in variables of separation
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{x}}\beta(z)=2\mathrm{a}\beta(z)-2\mathrm{b}\alpha(z),$
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{t}}\beta(z)=2(\mathrm{a}z+\alpha_{2N-2})\beta(z)-2(\mathrm{b}z+\beta_{2N-1})\alpha(z),$
where all dynamic variables are functions of $\mathrm{x}$ and $\mathrm{t}$. Therefore, zeros of $\beta(z)$ are functions of $\mathrm{x}$ and $\mathrm{t}$ as well, namely $\beta(z)=\mathrm{b}\prod_{k=1}^{N}(z-z_{k}(\mathrm{x},\mathrm{t}))$. Then
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{x}}\log\beta(z)=-\frac{1}{z-z_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}\mathrm{x}}=2\mathrm{a}-2\mathrm{b}\frac{\alpha(z)}{\beta(z)},\quad k=1,\dots,N,$
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{t}}\log\beta(z)=-\frac{1}{z-z_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}\mathrm{t}}=2(\mathrm{a}z+\alpha_{2N-2})-2(\mathrm{b}z+\beta_{2N-1})\frac{\alpha(z)}{\beta(z)}.$
@@
$\frac{\mathrm{d}z_{k}}{\mathrm{d}\mathrm{x}}=\frac{2w_{k}}{\prod_{j\neq k}^{N}(z_{k}-z_{j})},$ $\frac{\mathrm{d}z_{k}}{\mathrm{d}\mathrm{t}}=-\frac{2w_{k}\sum_{j\neq k}z_{j}}{\prod_{j\neq k}(z_{k}-z_{j})}.$
Now, let $D$ be a divisor of points $\{(z_{k},w_{k})\}_{k=1}^{N}$ defined by (36). The Abel image
$u=\mathcal{A}(D)=\sum_{k=1}^{N}\int_{\infty}^{(z_{k},w_{k})}\mathrm{d}u=\sum_{k=1}^{N}\int_{\infty}^{(z_{k},w_{k})}\begin{pmatrix}1\\ z\\ \vdots\\ z^{N-1}\end{pmatrix}\frac{\mathrm{b}\mathrm{d}z}{-2w}$
depends on $\mathrm{x}$ and $\mathrm{t}$, since the points $(z_{k},w_{k})$ are functions of $\mathrm{x}$ and $\mathrm{t}$. Then
$\frac{\mathrm{d}u_{2n-1}}{\mathrm{d}\mathrm{x}}=\sum_{k=1}^{N}\frac{\mathrm{b}z_{k}^{N-n}}{-2w_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}\mathrm{x}}=\sum_{k=1}^{N}\frac{-\mathrm{b}z_{k}^{N-n}}{\prod_{j\neq k}^{N}(z_{k}-z_{j})}=-\mathrm{b}\delta_{n,1},$
$\frac{\mathrm{d}u_{2n-1}}{\mathrm{d}\mathrm{t}}=\sum_{k=1}^{N}\frac{\mathrm{b}z_{k}^{N-n}}{-2w_{k}}\frac{\mathrm{d}z_{k}}{\mathrm{d}\mathrm{t}}=\sum_{k=1}^{N}\frac{\mathrm{b}z_{k}^{N-n}\sum_{j\neq k}z_{j}}{\prod_{j\neq k}(z_{k}-z_{j})}=-\mathrm{b}\delta_{n,2}.$
Thus, $u_{1}=-\mathrm{b}\mathrm{x}+C_{1}$, $u_{3}=-\mathrm{b}\mathrm{t}+C_{3}$, and $u_{2n-1}=C_{2n-1}=\operatorname{const}$, $n=3$, …, $N$.
@@
(48) $\beta(\mathrm{x},\mathrm{t})=-\mathrm{b}\wp_{1,1}(-\mathrm{b}\mathrm{x}+C_{1},-\mathrm{b}\mathrm{t}+C_{3},C_{5},\dots,C_{2N-1}).$
Since $\mathrm{b}^{2}\in\operatorname{\mathbb{R}}$, we have two possibilities: (i) $\mathrm{b}$ is real, or (ii) $\mathrm{b}$ is purely imaginary. In the case (i), the first two arguments $u_{1}$, $u_{3}$ of $\wp_{1,1}$ run along lines parallel to the real axes. In the case (ii), the first two arguments $u_{1}$, $u_{3}$ of $\wp_{1,1}$ go parallel to the imaginary axes. If $\beta$ describes a quasi-periodic wave, none of the mentioned lines coincides with the real or imaginary axis, due to the singularity of $\wp_{1,1}$ at $u=0$.
Next, we find the constant vector $\bm{C}=(C_{1},C_{3},\dots,C_{2N-1})$ such that $\beta$ in (48) is real-valued. We call this the reality conditions.